OpenStudy (anonymous):

Three ounces of cinnamon cost $2.40. If there are 16 ounces in 1 pound, how much does cinnamon cost per pound

OpenStudy (igreen):

2.4 = 1 ounce 2.4 * 16 = 1 pound 2.4 * 16 = 38.4 $38.40 = 1 pound

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @quanna15

OpenStudy (anonymous):

2.40/3 is 0.80 which is cinnamon per ounce, then 0.80*16 to get per pound 12.80$

OpenStudy (igreen):

Oops, didn't see that 2.4 = 3 ounce. 2/4 / 3 = 0.80. So 0.80 = 1 ounce. 0.80 * 16 = 1 pound 0.80 * 16 = 12.8 $12.80 = 1 pound